Solve this system of equations: -x+2y=11 and 3x-2y=-13
Please show work so I can understand :)

Answer by jim_thompson5910(35256)

Hint:
Solve for x in -x+2y=11 to get

-x+2y=11

-x=11-2y

x = -11+2y

x = 2y - 11

Then replace 'x' in 3x-2y = -13 with 2y - 11 to get

3x-2y = -13

3(2y-11)-2y = -13

I'll let you finish